In a ?ABC, ?C = p / 2, then 2 (r + R) is equal to

Grade:12

2 Answers

bharat bajaj IIT Delhi
10 years ago
?C = p / 2
R = c / (2 sin C) = c / 2

Now, 2 (r + R) = 2 r + 2 R
= 2 (s - c)tan(C / 2) + 2 (c / 2)
= 2s - 2c + c
= 2s - c
= a + b
Thanks
Bharat Bajaj
IIT Delhi
askiitians faculty
bharat bajaj IIT Delhi
10 years ago
?C = p / 2
R = c / (2 sin C) = c / 2

Now, 2 (r + R) = 2 r + 2 R
= 2 (s - c)tan(C / 2) + 2 (c / 2)
= 2s - 2c + c
= 2s - c
= a + b
Thanks
Bharat Bajaj
IIT Delhi
askiitians faculty